OpenStudy (anonymous):

derivative of tan^2x+sinx

OpenStudy (anonymous):

i rewrote the function as (tanx)^2+sinx

OpenStudy (anonymous):

=tan(x)tan(x) + sin(x) you need to use the product rule on the first part, and the second part is simple derivation so let u=tan(x) v=tan(x) so the derivative of tan(^2)x will be uv'+vu' then add the derivative of sin(x) on the end does that make sense? if you struggle to differentiate tan(x), think of it as sinx/cosx and use the quotient rule

OpenStudy (anonymous):

okay so i have f'(x)= 2tanxsec^2x+cosx

OpenStudy (anonymous):

yep that's correct! :)

OpenStudy (anonymous):

the question asks me to do f'(pi/4) how can i plug that in?

OpenStudy (anonymous):

okay so if you know f '(x) ie what you've just worked out, simply replace all of your x's by pi/4 and you'll get your answer!

OpenStudy (anonymous):

yes but how can i plug it in into sec^2x? my calculator doesnt let me do that

OpenStudy (anonymous):

oh okay i see what you mean. you can just do 1/[(cos x)*cos(x)]

OpenStudy (anonymous):

ooh okay got it thank you:)

OpenStudy (anonymous):

you're welcome! good work :)